+3 7 اذا كانت الدالة - 22( z2 + 2) = (2)f فان (0,Res ( f هو ه م 2 م لا | -
Answer #1

Problem given that Zt3 f(2)= 12+2) : ZZ(2²+2) 2=0 has a pole of order 2 then Res (4,0) = lim (2242) Res (4,0) = lim dl zoo al